OpenStudy (anonymous):

Solve x2 + 4x − 4 = 0. x equals the quantity of negative 2 plus or minus square root 16 all over 2 x equals negative 4 plus or minus square root 2 x equals negative 2 plus or minus 2 square root 2 x equals the quantity of 4 plus or minus square root 18 all over 2

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@dan815

Nnesha (nnesha):

use the quadratic formula

Nnesha (nnesha):

\[\huge\rm \frac{ -b \pm \sqrt{b2-4ac} }{ 2a }\] a ,b,c are \[\rm Ax^2+Bx+C=0\] a=leading coefficient b=middle term c=constant plugin

OpenStudy (anonymous):

x² + 4x + 4 = 4 + 4 (x + 2)² = 8 √(x + 2)² = ±√8 x + 2 = ±2√2 x = -2 ± 2√2

Nnesha (nnesha):

looks right .......

OpenStudy (anonymous):

ok